### v2

# Anatomical feature-prioritized loss for enhanced MR to CT translation

Abstract
In medical image synthesis, the precision of localized structural details is crucial, particularly when addressing specific clinical requirements such as the identification and measurement of fine structures. Traditional methods for image translation and synthesis are generally optimized for global image reconstruction but often fall short in providing the finesse required for detailed local analysis. This study represents a step toward addressing this challenge by introducing a novel anatomical feature-prioritized (AFP) loss function into the synthesis process. This method enhances the reconstruction by focusing on clinically significant structures, utilizing features from a pre-trained model designed for a specific downstream task, such as the segmentation of particular anatomical regions. The AFP loss function can be used to replace or complement global reconstruction methods, ensuring a balanced emphasis on both global image fidelity and local structural details. Various implementations of this loss function are explored, including its integration into different synthesis networks such as GAN-based model and CNN-based model. Our approach is applied and evaluated in two contexts: lung MR to CT translation, focusing on high-quality reconstruction of bronchial structures, using a private dataset; and pelvis MR to CT synthesis, targeting the accurate representation of organs and muscles, utilizing a public dataset from the Synthrad2023 challenge. We leverage embeddings from pre-trained segmentation models specific to these anatomical regions to demonstrate the capability of the AFP loss to prioritize and accurately reconstruct essential features. In the context of lung airway segmentation, the Dice coefficient was enhanced from 0.5342 using L1 loss to 0.5841 with the adoption of AFP loss. Similarly, in pelvic applications, the Dice coefficient for bone reconstruction in synthesized CTs improved from 0.7381 when employing L1 loss to 0.7800 with AFP loss. This tailored approach shows promising potential for enhancing the specificity and practicality of medical image synthesis in clinical applications.
Keywords Cross-Modality Translation, Fine-Structure Preservation, Lung MRI
1 Introduction
Over the past decade, medical image synthesis has emerged as a significant trend in the field of healthcare. This technique plays a pivotal role in producing synthetic data to address the shortfall in medical datasets, such as data diversity ([1]), and improving the robustness of machine learning models. Additionally, it serves as a valuable tool for translating between different imaging modalities, offering benefits in applications like attenuation correction or radiation therapy planning, as described in [2]. The significant advancement of this field is mainly attributed to the evolution of deep learning, notably the advent of generative methods, such as generative adversarial networks (GANs) ([3]) and diffusion models (DMs) ([4]). Many recent studies have explored synthetic data generation ([5]) as well as cross-modality translation, addressing reconstruction challenges in anatomical areas such as brain ([6]), aorta ([7]), lungs ([8]).
Despite these advances, there remains a gap in the precision and fidelity of synthesized images, particularly in the reconstruction of detailed anatomical structures. Most current state-of-the-art methods, including those based on conditional generative models (e.g. cGANs introduced in [9]) focus on optimizing global image reconstruction, often at the expense of accurately capturing the nuances of specific features. This is largely due to the reliance on global loss functions such as the L1 loss, which primarily measures global intensity differences and does not account for the preservation of anatomical structures and critical medical features. This kind of loss also tends to smooth out finer details, leading to images that may appear realistic at a macro level but lack the precision required for clinical tasks involving the identification and precise measurement of small, intricate structures. This oversight becomes particularly critical when addressing explicit clinical needs, such as the accurate delineation of boundaries between different tissues, the identification of lesions, or the precise measurement of anatomical structures for surgical planning or disease monitoring.
Several papers have focused on cross-modality synthesis based on anatomical structures. In [10], authors use a structure-constrained GAN for unsupervised MR-to-CT synthesis by defining a structure-consistency loss based on the modality independent neighborhood descriptor, introduced in [11], often employed for cross-modality registration. [12] employs a multi-task network for synthetic CT generation from MRI, emphasizing bone density value prediction. It uses a composite loss function to localize regions of interest, combining classification and regression tasks to improve performance and computational efficiency, particularly in radiation therapy planning. [13] introduces a structure-aware GAN for organ-preserving synthetic CT generation from MRI, using a dual-stream approach based on a segmentation loss to maintain organ integrity and achieve clinically acceptable accuracy in MR-only treatment planning.
These methods focus on specific body parts or anatomical structures using segmentation networks, but they do not use the full capabilities of these networks. While employing a segmentation loss is interesting for cross-modality synthesis, it typically relies only on the activation of the last layer of the network and is limited by the specific labels used, neglecting the rich representations available across various network layers. Addressing these limitations, our study introduces a novel approach aimed at enhancing the synthesis of specific medical details across different anatomical regions. We introduce an anatomical feature-prioritized loss function, denoted AFP loss, specifically designed for medical image translation, ensuring both global fidelity and detailed structural accuracy. This approach leverages a distance between features extracted from a model pretrained on a specific medical task, such as the segmentation of a delimited region, based on the idea that such a model presents an advanced understanding and representation of those medical structures.
We integrate this feature-prioritized loss function into various synthesis networks, including generative models like GANs (pix2pixHD from [14], SPADE from [15]) and adapted CNN models (nnU-Net from [16]), to evaluate their effectiveness comprehensively. We then assess the performance of our approach on two distinct tasks: MR to CT synthesis in the lungs, aiming for fine structure generation including vessels and bronchi, and MR to CT synthesis in the pelvis, focusing on the generation of coarser anatomical features such as organs and muscles. Our study provides a robust qualitative and quantitative evaluation, focusing on the precise reconstruction of anatomical structures in these regions, demonstrating the potential of our approach to improve clinical outcomes.
2 Related works
2.1 Image Synthesis Models
CNN-based Models
Convolutional Neural Networks (CNNs) like U-Net ([17]) have been foundational in medical image synthesis due to their ability to capture both local and global features efficiently. U-Net’s encoder-decoder architecture with skip connections is particularly effective for tasks requiring precise localization, such as segmentation and synthesis.
Generative Adversarial Networks (GANs)
GANs and their conditional variants (cGANs) like pix2pix and cycleGAN have significantly advanced the field by enabling high-quality image-to-image translation. Pix2pix ([18]) uses paired data to learn the mapping from input to output images, while cycleGAN ([19]) enables translation between unpaired datasets, making it highly versatile. [15] introduces SPADE, which further improves semantic image synthesis by incorporating a spatially-adaptive normalization.
Diffusion Models
Diffusion models, particularly those employing latent variable frameworks, have shown promise in generating high-resolution images. Latent Diffusion Models ([4]) offer a robust alternative by focusing on iterative refinement of image quality through a stochastic process, outperforming traditional GANs in several benchmarks.
2.2 Insights from Segmentation Tasks
The nnU-Net framework is kwown as the state-of-the-art model in medical image segmentation, demonstrating superior performance across a wide range of datasets and challenges ([16]). Its automated configuration of network architectures and hyperparameters ensures optimal performance tailored to specific datasets, making it highly adaptable and efficient. Recent studies have shown that nnU-Net consistently outperforms more complex architectures, including transformer-based models, in terms of segmentation accuracy and robustness. This highlights the effectiveness of well-tuned CNN architectures in medical imaging tasks.
The potential of nnU-Net for translation tasks has been demonstrated in [20], where it matches or outperforms GAN-based models in cross-modality synthesis. The main advantage of using this framework is that it allows us to redefine the loss functions and not be limited by the losses of GANs, which are often based on visual fidelity and are prone to divergence. Relying on a simple, effective model like nnU-Net allows us to highlight the effectiveness of our AFP loss, either as a standalone or coupled with an L1 loss.
3 Method
3.1 Anatomical Feature-Prioritized Loss Function
fig:layers Refer to caption
Figure 1: Example of an MR to CT translation pipeline using a feature-prioritized loss on a U-Net network trained for airways segmentation. All features are acquired after the ReLU activation in each convolutional block. Layer 7 is extracted before the final convolution.
In this study, we introduce a novel anatomical feature-prioritized loss function (denoted AFP loss). The primary goal of this loss is to enhance the synthesis of specific anatomical structures in medical images, which are often inadequately reconstructed when relying solely on standard reconstruction losses. Our approach leverages a pre-trained segmentation convolutional network, specifically tailored to segment precise anatomical regions. The specialization of this network in segmenting precise regions allows us to leverage its data representations to improve the generation of anatomical features. The mechanism of our feature-prioritized loss mirrors that of a perceptual loss described in [21], primarily focusing on minimizing the difference between the feature representations of synthetic and real images extracted from various layers of a CNN. Traditionally, perceptual loss is defined using 2D classification networks on natural images, such as VGG on ImageNet, and its use in 3D is often limited. Our anatomical feature-prioritized loss, in contrast, leverages segmentation networks, making it more straightforward to implement in medical tasks. Additionally, it benefits from more precise feature representations compared to those in classification models. The AFP loss is formulated as follows :
$\mathcal{L}_{AFP(x,y)}=\frac{1}{N}\sum_{i=1}^{N}\left\|\phi_{i}(x)-\phi_{i}(y)\right\|_{1}$ (1)
Here, $x$ and $y$ represent the synthesized and real images to compare, $N$ denotes the number of layers extracted from the pre-trained network and $\phi_{i}$ the features of the i-th layer.
3.2 Integration in synthesis networks
The AFP loss introduced in this study is a versatile function, adaptable to various synthesis and translation tasks within medical imaging. Depending on the specific synthesis task, the AFP loss can be combined with a global reconstruction loss, such as the traditional L1 loss, to maintain a correct range of intensities. A general recommendation is to start with a global network trained with L1 loss to establish a baseline intensity reconstruction and then fine-tune the model with our AFP loss to refine the anatomical structures. This approach ensures a two-step process: an initial global generation step followed by a local anatomical refinement step.
For tasks involving the generation of fine structures, such as vessels or bronchi, it may be preferable to use only the AFP loss from the beginning. In such cases, L1 loss should be avoided due to its tendency to produce blurred reconstructions, which can obscure fine details. The AFP loss excels in these scenarios by focusing on precise anatomical fidelity without being influenced by global intensity differences.
While we illustrate this loss using pre-trained U-Net models, our approach can be adapted to any CNN trained on medical data. The AFP loss serves as a robust addition to medical image synthesis and translation tasks, enhancing the capability to generate high-fidelity anatomical structures.
3.3 Implemented Synthesis Models
To evaluate the performances of our anatomical structure prioritized loss function (AFP loss), we implemented it in two primary models for medical image synthesis: a GAN-based model and a CNN-based models. Specifically, we develop a 3D variant of the pix2pixHD model and adapt a 3D nnU-Net model for translation tasks.
GAN-based model
We develop a 3D variant of the pix2pixHD model ([14]), incorporating 3D versions of the generator and discriminator, as well as 3D SPADE layers ([15]). This architecture facilitates image-to-image translation by incorporating spatially-adaptive normalization, effectively mitigating the issue of false positives encountered in the pix2pixHD network alone ([22]). For the training process, we employ a composite loss function that includes an adversarial loss, a feature-matching loss, and a global reconstruction loss. The adversarial loss utilizes a Hinge loss mechanism, as adopted in the SPADE implementation ([15]), to refine the adversarial training dynamics. The feature-matching loss ensures alignment between the discriminator’s intermediate representations of the real and synthesized images, utilizing an L1 distance to promote stability and fidelity in the generated images ([14]). A global reconstruction loss is incorporated to further enhance the overall quality of the synthesized output, defined as an L1 loss. Finally, our ASP loss is incorporated to complement the L1 loss, enhancing localized structural details, or to replace it entirely, offering a more targeted approach to preserving anatomical accuracy in the synthesized images.
CNN-based model
We have extended the application of the nnU-Net framework ([16]), traditionally a pioneer in medical image segmentation, to tackle translation tasks. Our previous work ([20]) provides an adaptation of the nnU-Net for synthesis tasks and showed promise with a standard MSE loss function. The adapted model retains the original framework’s preprocessing and low-complexity architecture, which benefits from not being dependent on a discriminator, a component known for its training challenges. We propose a new version of the adapted nnU-Net incorporating the proposed AFP loss in the training process. This setup enables us to precisely quantify the impact of our AFP loss, whether used alone or paired with an L1 loss, providing a robust framework for improving medical image synthesis. The model has been made publicly available at https://github.com/Phyrise/nnU-Net_translation.
3.4 Reconstruction from patches
To address the memory constraints associated with processing 3D models, we adopt a patch-based approach for each synthesis model. To ensure smooth reconstruction and mitigate border effects between patches, we utilize standard median reconstruction on voxel intensities, with patches defined by a tiling of 0.5. This method preserves high-frequency details, resulting in a visually more accurate and detailed final output. While mean reconstruction often acts as a smoothing filter that can perform better on certain metrics, median reconstruction provides superior visual quality by maintaining the finer details.
3.5 Checkerboard artifacts
The AFP loss operates similarly to a perceptual loss, which has been shown to potentially lead to checkerboard artifacts ([23]), primarly due to transposed convolution operations in the decoder part. To adress this issue, we explore the effectiveness of combining the AFP loss with a traditional L1 loss, which can be effective in reducing high-frequency errors that contribute to checkerboard patterns. This approach aim to leverage the global intensity preservation of the L1 loss while benefiting from the detailed structure preservation of the AFP loss. We also investigated replacing transposed convolutions with a combination of upsampling and convolutional layers as proposed by [24]. This proven technique helps to reduce checkerboard artifacts by providing a more stable upsampling mechanism, particularly when using feature-based losses.
4 Experiments
4.1 Datasets
Lung MR to CT Synthesis
For lung MR to CT synthesis, we used a private dataset of thoracic images from UTE MR and CT scans of 122 patients. Both modalities were acquired on the same day, from 2018 to 2022. CT images were obtained using a Siemens SOMATOM Force and a GE Medical Systems Revolution CT in end-expiration, with sharp filters. The acquisition parameters were a DLP of 10 mGy.cm and a SAFIRE iterative reconstruction. The UTE MR images were acquired using the SpiralVibe sequence on a SIEMENS Aera scanner, with the following parameters: TR/TE/flip angle=4.1ms/0.07ms/5°. Since the slice plane is encoded in Cartesian mode, the native acquisition was performed in the coronal plane with field-of-view outside the anterior and posterior chest edges to prevent aliasing. It should be noted that resolutions, voxel spacings, and fields of view are not identical in CT and MR images. In addition, modalities may have been taken at different points in the respiratory cycle.
All volumes are resampled to have a voxel size of 0.6 $\times$ 0.6 $\times$ 0.6 mm3. To align CT images with MR images, we first perform a rigid translation, followed by a deformable registration process using the EVolution algorithm proposed by [25]. This method incorporates a similarity term that prioritizes the alignment of edges and employs a diffeomorphic transformation to preserve the CT volume topology. Subsequently, we normalize MR images to have zero mean and unit variance, and for CT images, we use intensity values from the foreground classes to compute the mean and standard deviation, clip values to the 0.5 and 99.5 percentiles, and then normalize by subtracting the mean and dividing by the standard deviation.
Pelvis MR to CT Synthesis
For the synthesis of pelvis images, we used the open-access dataset from Task 1 of the SynthRAD2023 challenge from [26]. This dataset includes paired MR and CT images from 180 patients used for training, with an additional 30 patients designated for validation. The dataset provides high-quality, well-aligned MR and CT image pairs, which eliminates the need for additional registration processes. The voxel sizes in this dataset are already standardized to 2.5 × 1 × 1 mm, ensuring consistent spatial resolution across all images. Similar preprocessing steps from the lung dataset are applied, with z-score normalization for MR images and clipping CT values to the 0.5 and 99.5 percentiles, followed by normalization using the mean and standard deviation.
4.2 Evaluation Metrics
To comprehensively assess our synthesis models, we employ a dual approach combining intensity-based and task-specific metrics.
Intensity-based Metrics
Traditional intensity-based metrics, such as Mean Absolute Error (MAE) and Structural Similarity Index (SSIM), provide a measure of global image synthesis quality but may not adequately capture the nuances required for specific medical tasks. Accurate delineation of boundaries between different tissues, identification of lesions, and precise measurement of anatomical structures for surgical planning or disease monitoring necessitate evaluation metrics based on the reconstruction of anatomical structures. Although qualitative assessments by expert radiologists are valuable, they are also very tedious and time-consuming.
Task-specific Metrics
To capture the nuances required for specific medical tasks, we introduce a downstream task - the segmentation of anatomical structures - as a key evaluation method. This approach provides a more precise measure of the model’s performance at the anatomical level and helps quantify the false positive rate, crucial for clinical applications. We use segmentation-based metrics like the Dice score and the Normalized Surface Distance (NSD) ([27]) to evaluate local reconstructions. NSD is particularly valuable as it includes a manually defined margin of error, enhancing its utility in analyzing fine structures and providing robustness against registration inaccuracies. In our study, the margin of error for NSD calculations is set at twice the voxel size: 1.2mm for lungs and 2mm for pelvis. To mitigate potential biases introduced by relying on pretrained segmentation networks, we evaluate our synthesis models across multiple segmentation tasks and models. This ensures a comprehensive and balanced assessment of their performance. For both tasks, the segmentation of the real CT is defined as a silver-standard, given that the actual segmentation masks are not available.
Lung MR to CT Synthesis
For the lung region, we use segmentations from three pre-trained models:
  1. 1.
    The public NaviAirway network ([28]), focused on segmenting airways and bronchi.
  2. 2.
    The publicly available TotalSegmentator ([29]) models, which have been fine-tuned and fused into one comprehensive model. This network focuses on coarser segmentations such as lung lobes, bones, muscles, and organs outside the lungs.
Pelvis MR to CT Synthesis
For the pelvis region, we use segmentations from the TotalSegmentator model. This network provides segmentation of various tissues in the pelvic region, including bones (e.g., sacrum, hips), organs (e.g., bladder, prostate), and muscles (e.g., iliopsoas, gluteus).
This dual approach of intensity-based and task-specific metrics allows for a comprehensive evaluation of both global image quality and specific anatomical structure preservation in our synthesis models.
Refer to caption Refer to caption Refer to caption
Input MR Real CT L1
Refer to caption Refer to caption Refer to caption
MedicalNet AFP loss w/ AFP loss w/
TotalSegmentator NaviAirway
Figure 2: Comparison between input MR image, ground truth CT, and synthesized CT from adapted nnU-Net with several loss functions. Coronal slices are presented, including a zoomed-in view of the bronchi reconstruction. The zoomed-in view demonstrates that the AFP loss with NaviAirway features enables the most faithful reconstruction of the real CT.
Table 1: Comparison of 3D lung networks performances using MAE and SSIM between real CT and synthesized CT, as well as Dice score and NSD on airways segmentation using the NaviAirway pipeline. Results include adapted nnU-Net model with various losses and GAN-based SPADE model.
Intensity-based metrics Task-based metrics
Model / Loss MAE SSIM Dice NSD
nnU-Net / L1 $\mathbf{48.7198\pm 12.8441}$ $\mathbf{0.8367\pm 0.0152}$ $0.5342\pm 0.0655$ $0.6442\pm 0.0898$
nnU-Net / MedicalNet $69.6625\pm 19.7463$ $0.8146\pm 0.0165$ $0.4711\pm 0.0648$ $0.5907\pm 0.0742$
nnU-Net / AFP TotalSeg $59.3991\pm 14.9090$ $0.8287\pm 0.0146$ $0.5307\pm 0.0682$ $0.6674\pm 0.0764$
nnU-Net / AFP Navi $64.9802\pm 16.8825$ $0.8243\pm 0.0144$ $0.5801\pm 0.0592$ $0.7206\pm 0.0637$
SPADE / L1 $55.6784\pm 14.4213$ $0.8298\pm 0.0158$ $0.5212\pm 0.0627$ $0.6238\pm 0.0857$
SPADE / AFP Navi $69.1460\pm 18.2897$ $0.8155\pm 0.0167$ $0.5459\pm 0.0712$ $0.6891\pm 0.0791$
5 Results
5.1 Lung MR to CT Synthesis
Refer to caption Refer to caption Refer to caption
Real CT L1 AFP Navi
Figure 3: Example of 3D airway segmentation using the NaviAirway pipeline. The trachea, in transparency, is excluded for the calculation of metrics. The AFP loss facilitates the reconstruction of higher bronchi compared to the L1 loss, as depicted in Table 1.
Table 2: Comparison of 3D nnU-Net models with different losses for lung synthesis, evaluated using Dice score and NSD on TotalSegmentator segmentations between real CT and synthesized CT.
Loss Muscle/Organs Bones Lobes
Dice NSD Dice NSD Dice NSD
L1 $0.7937\pm 0.0526$ $0.8712\pm 0.0580$ $0.4015\pm 0.0808$ $0.4933\pm 0.0995$ $0.8560\pm 0.0704$ $0.8897\pm 0.0727$
MedicalNet $0.7737\pm 0.0434$ $0.8501\pm 0.0476$ $0.4051\pm 0.0578$ $0.4987\pm 0.0710$ $0.8811\pm 0.0760$ $0.9095\pm 0.0790$
AFP TotalSeg $\mathbf{0.8107\pm 0.0568}$ $\mathbf{0.8999\pm 0.0628}$ $\mathbf{0.5157\pm 0.0699}$ $\mathbf{0.6292\pm 0.0864}$ $\mathbf{0.9264\pm 0.0466}$ $\mathbf{0.9598\pm 0.0483}$
AFP Navi $0.7509\pm 0.0628$ $0.8010\pm 0.0807$ $0.4785\pm 0.0735$ $0.5827\pm 0.0917$ $0.8601\pm 0.0803$ $0.8915\pm 0.0835$
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Real CT L1 MedicalNet AFP TotalSeg AFP NaviAirway
Figure 4: Comparison of lung labels segmentations from real CT and synthesized CTs using the TotalSegmentator pipeline. Bones are represented in yellow, lung lobes in blue, green, brown and red colors. Visual results align with Table 2 : The FP loss achieves superior reconstruction, particularly in the bones.
In the lung MR to CT synthesis task, we established the baseline using the traditional L1 loss and a perceptual loss leveraging MedicalNet’s embeddings. We evaluated these losses against our proposed AFP loss, using various pre-trained segmentation networks: TotalSegmentator (TotalSeg), NaviAirway (Navi). Fig. 2 presents a visual comparison of lung coronal sections from different models. The model using L1 loss results in blurred images, with particularly poor bone reconstruction. In contrast, the MedicalNet-based model produces sharper images but introduces artifacts in anatomical areas. Models based on AFP loss offer visually appealing results, with well-reconstructed anatomical bones and lung tissue. Additionally, only NaviAirway provide detailed bronchial reconstruction, as highlighted in the zoomed-in section.
Table 1 presents a quantitative evaluation of the model’s performance on lung MR to CT synthesis, based on the MAE, SSIM, Dice score, and NSD between synthesized and ground truth CT images. The adapted nnU-Net trained with L1 loss delivers the best performance on intensity-based metrics. Other models based on perceptual loss or AFP loss achieve average results in terms of MAE but maintain reasonable SSIM values. In the context of airway segmentations using the NaviAirway pipeline, the adapted nnU-Net models with AFP loss from NaviAirway’s embeddings deliver the best performance, with the highest Dice scores and NSD values. Conversely, models employing L1, perceptual, or AFP loss with TotalSegmentator’s embeddings yield poorer results, lacking precise bronchial reconstruction. The GAN-based SPADE method generally underperforms compared to the adapted nnU-Net, but adding AFP loss to SPADE enhances its performance. These metrics align with qualitative analysis from Fig. 2, with the models using AFP loss delivering the best performance in airway reconstruction.
Table 2 and Fig. 4 present a comparison of the models on the reconstruction of lung anatomical regions using the TotalSegmentator (TotalSeg) pipeline with Dice score and NSD. The model trained with the AFP loss using TotalSeg’s embeddings provide the best performances. A combination of AFP loss using NaviAirway’s embeddings also delivers positive results, whereas other models exhibit notably poorer reconstructions especially in the bones.
Table 3: Comparison of 3D nnU-Net models with different losses for pelvic synthesis using Dice score and NSD on TotalSegmentator segmentations between real CT and synthesized CT.
Loss Bones Muscles Organs
Dice NSD Dice NSD Dice NSD
L1 $0.7381\pm 0.2394$ $0.7743\pm 0.2217$ $0.6869\pm 0.2101$ $0.7189\pm 0.2081$ $0.7636\pm 0.2272$ $0.7943\pm 0.2073$
MedicalNet $0.7274\pm 0.2411$ $0.7585\pm 0.2361$ $0.6462\pm 0.2373$ $0.6983\pm 0.2254$ $0.7219\pm 0.2359$ $0.7684\pm 0.2178$
L1 + AFP TotalSeg $0.7716\pm 0.1942$ $0.8173\pm 0.1957$ $\mathbf{0.7281\pm 0.1947}$ $\mathbf{0.7683\pm 0.1848}$ $0.8004\pm 0.1979$ $0.8362\pm 0.1837$
AFP TotalSeg $\mathbf{0.7800\pm 0.2030}$ $\mathbf{0.8227\pm 0.2001}$ $0.7240\pm 0.1927$ $0.7617\pm 0.1884$ $\mathbf{0.8061\pm 0.1877}$ $\mathbf{0.8482\pm 0.1784}$
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Input MR Real CT L1 MedicalNet AFP TotalSeg L1 + AFP TotalSeg
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5: Comparison between input MR image, ground truth CT, and synthesized CT from nnU-Net translation with several loss functions. Sagittal slices are presented, including a zoomed-in view of the colon reconstruction on the top row, and 3D view of TotalSegmentator segmentations on the bottom row. Bones are represented in yellow, muscle in brown, organs in blue and colon region in green. Visual results align with Table 3 and show that AFP loss enables a better anatomical reconstruction, especially in the organs and colon.
5.2 Pelvis MR to CT Synthesis
Similarly to the lung task, in the pelvic image synthesis task, we established a baseline using the traditional L1 loss and perceptual loss using the MedicalNet model. Our AFP loss, leveraging embeddings from the TotalSegmentator (denoted TotalSeg) network, was evaluated against this baseline. Furthermore, we explored a hybrid approach that combines L1 loss with AFP loss (denoted L1 + AFP) to assess the benefits of retaining the global reconstruction L1 loss.
Table 4 presents a quantitative evaluation of the model’s performance on pelvic MR to CT synthesis, based on the MAE and SSIM between synthesized and ground truth CT images. Similarly to the lung results, the model trained with L1 loss delivers the best performance on intensity-based metrics. Other models based on perceptual loss or AFP loss alone achieve average results in terms of MAE but maintain reasonable SSIM values. Finally, the combination of L1 and AFP losses provides solid results on these intensity-based metrics, bringing it closer to the high standards set by L1 loss.
Table 4: Comparison of 3D nnU-Net models with different losses for pelvic synthesis using MAE and SSIM between real CT and synthesized CT.
Loss MAE SSIM
L1 $\mathbf{65.3000\pm 15.1801}$ $\mathbf{0.8520\pm 0.0354}$
MedicalNet $81.3590\pm 30.7177$ $0.83852\pm 0.0383$
L1 + AFP TotalSeg $66.9377\pm 15.1169$ $0.8471\pm 0.0357$
AFP TotalSeg $85.2581\pm 14.4719$ $0.8390\pm 0.0366$
Fig. 5 presents a visual comparison of pelvic sagittal sections from different models. The model trained with L1 loss displays blurry results, with smooth intensities across the body and a noticeable lack of precise structural delineations. Conversely, the MedicalNet-based model, while producing sharper images, tends to introduce artifacts and also falls short in defining precise anatomical structure borders. In contrast, the TotalSegmentator model delivers excellent visual quality, featuring sharp reconstruction and precise delineation of structures, notably in the colon region as depicted in the zoomed-in preview.
Table 3 and bottom row of Fig. 5 provide a comparative evaluation of the models on anatomical regions using the TotalSegmentator pipeline with Dice score and NSD. Aligning with the visual analysis, the model trained with the AFP loss using TotalSeg’s embeddings provide the best performances in every region. Combining the AFP loss with a L1 loss does not significantly affect the reconstruction performance of anatomical structures. The model trained with MedicalNet perceptual loss provides poor results overall, while the model trained with L1 loss delivers average performance in each region.
6 Discussion
Feature-Prioritized Loss: A Simple Yet Effective Approach
This study presents an exploration into medical image synthesis with a focus on MR to CT translation for both lung and pelvic regions, specifically aiming to enhance localized structural detail reconstruction through a novel anatomical feature-prioritized loss (AFP loss). Our findings reveal that this method effectively reconstructs targeted anatomical features and can either complement or replace global reconstruction strategies such as L1 loss or perceptual loss. The AFP loss, while conceptually simple in its implementation—essentially applying L1 loss in the embeddings of a pretrained segmentation network—demonstrates a significant impact on the representation of anatomical structures. This straightforward approach allows the synthesis model to develop a more advanced understanding and representation of complex medical structures, highlighting the potential of leveraging domain-specific knowledge in synthesis networks.
Comparison with Traditional Methods and Metrics
Global reconstruction loss is commonly expressed using a perceptual loss, an L1 loss, or an L2 loss. The standard L1 loss, directly applied to voxel intensities, is usually defined as the primary global reconstruction loss, as it is more robust to outliers in contrast to the L2 loss, which also tends to yield blurry results ([18]). On the other hand, perceptual losses are often more robust and deliver better results than L1 loss but are usually limited to 2D applications [30]), unlike our AFP loss, which can easily be used in 3D using trained segmentation networks. However, the effectiveness of the feature-prioritized loss is contingent upon the performance of the pretrained model on the given task, as it relies on the auxiliary model’s ability to provide meaningful gradients for the synthesis model’s training. Our experiments showed that the AFP loss could significantly enhance the reconstruction quality of specific anatomical structures when integrated into various synthesis networks, including GAN-based models (pix2pixHD, SPADE) and CNN-based models (adapted nnU-Net).
For global image synthesis quality, traditional metrics such as MAE and SSIM were employed. However, these intensity-based metrics alone often do not adequately capture the nuanced needs of medical imaging, such as the accurate delineation of boundaries between different tissues or the identification of lesions. This underscores the necessity of using structure-specific metrics such as Dice score and Normalized Surface Distance (NSD) for evaluating the quality of local reconstructions, particularly focusing on preserving anatomical structural details that are vital for clinical analysis.
Performance and Insights
Our experiments indicate that the AFP loss excels in refining the reconstruction of targeted structures, regardless of the dataset (lung or pelvis) or the method used (GAN or U-Net). In the lung region, we evaluated our models on three segmentation tasks: airways and sain bronchi using the NaviAirway pipeline, and larger anatomical regions (muscles, organs, bones) using the TotalSegmentator pipeline. As expected, the best model for each task is often the one with embeddings closely aligned with the task being evaluated. For instance, for muscle and organ reconstruction based on TotalSegmentator segmentation, the most effective model utilizes AFP loss with TotalSegmentator’s embeddings.
However, surprisingly, some networks with non-specialized embeddings can produce high-quality reconstructions or even outperform specialized models when combined with them. For bone reconstruction, AFP loss using NaviAirway’s embeddings yielded excellent results, despite the network originally being specialized in fine structures within the lungs, which are quite different from bone segmentation. These findings highlight the advanced understanding and representation capabilities of pre-trained segmentation networks and underscore the value of using them as replacements for perceptual loss or L1 loss.
Moreover, while a general model like TotalSegmentator is not perfectly suited for fine structure reconstruction, it serves as an excellent starting point for coarser tasks. It systematically outperforms L1 loss and MedicalNet perceptual loss in every segmentation task, and provides sharper and more realistic visual reconstructions compared to these other models.
In the pelvic region, models using AFP loss on TotalSegmentator’s embeddings and AFP loss combined with L1 loss performed similarly in the downstream task and had similar visual reconstructions, despite having very different results in the intensity-based evaluation. This discrepancy once again calls into question the relevance of these metrics, as they do not accurately reflect the medical and anatomical accuracy of the synthesized volumes.
Initially, perceptual loss, typically derived from pretrained 2D networks like VGG, was considered. However, when applied using a 3D MedicalNet model specifically adapted for medical data, the results were underwhelming, even when compared to the traditional L1 loss. This suggests ongoing challenges in using medical image datasets for perceptual loss due to their limited diversity and smaller scale compared to extensive, varied datasets like ImageNet. Consequently, we recommand relying on the L1 loss or AFP loss with TotalSegmentator’s embeddings for coarse reconstructions, which consistently delivered superior image quality.
Moreover, the elimination of the discriminator, particularly when using nnU-Net as a generator only, resulted in even better outcomes. This highlights the potential need to reconsider traditional cost functions in medical synthesis. Instead of focusing on output likelihood — which is typical with discriminators — a more clinically relevant approach would focus on preserving anatomical structures and ensuring low false-positive rates. This aligns directly with the clinical objective of synthesizing volumes that not only possess high signal quality but also maintain the integrity of anatomical details, facilitating advanced analysis by radiologists.
7 Conclusion
Our study introduces an innovative approach to medical image synthesis, centered on an Anatomical Feature-Prioritized (AFP) loss for precise reconstruction of anatomical regions. The AFP loss, while conceptually simple—essentially applying L1 loss in the embeddings of a pretrained segmentation network—demonstrates a significant impact on the representation of anatomical structures. This straightforward approach allows the synthesis model to develop a more advanced understanding and representation of complex medical structures, highlighting the potential of leveraging domain-specific knowledge in synthesis networks. We explored strategies with this method in both lung and pelvis MR to CT synthesis using 3D SPADE and adapted nnU-Net models, demonstrating improved performance in both overall image quality and the intricate reconstruction of anatomical structures on three downstream tasks. Our results highlight the adaptability of the AFP loss in enhancing localized structural detail across different datasets and synthesis methods. Interestingly, our findings indicate that using AFP loss alone, particularly without a discriminator, can yield superior results in structure reconstruction, thus redefining cost functions in medical synthesis. A key aspect of our approach is the dependency on robust segmentation models. As long as a robust segmentation model is available, the loss can be computed effectively for the translation task, ensuring precise reconstruction of anatomical structures. Future work will extend the utility of our feature-prioritized loss across various segmentation networks and medical regions, and explore dynamic weighting approaches for each feature map layer. In conclusion, the AFP loss presents a promising direction for enhancing the precision of medical image synthesis, aligning more closely with clinical objectives by ensuring high fidelity in anatomical detail and reducing false positives. This approach holds significant potential for advancing the field and improving outcomes in medical imaging applications.